# A search for an eV-scale sterile neutrino using improved high-energy $\nu_{\mu}$ event reconstruction in IceCube

### v1

@@ Abstract
Stringent constraints on sterile-neutrino-induced $\nu_{\mu}$ disappearance arise from IceCube’s study neutrinos produced in the atmosphere with TeV energies, known as high-energy atmospheric neutrinos [18, 19, 20]. IceCube, a gigaton ice-Cherenkov detector near the geographic South Pole [21], detects neutrinos produced in cosmic-ray air showers. The detector consists of an array of photomultiplier tubes in glass pressure housings called “Digital Optical Modules” (DOMs) [22], deployed in 86 strings. In the presence of a sterile neutrino, high-energy atmospheric neutrinos detected by IceCube experience oscillation effects beyond vacuum-like oscillations [23, 24, 25, 26, 27] including, for eV-scale masses, matter-enhanced resonant effects induced as high-energy antineutrinos cross the Earth’s core [28, 29, 30, 31, 32, 33]. Notably, while standard three-neutrino oscillations are negligible at TeV energies due to their mass splitting, the presence of an eV-scale sterile neutrino yields distinct $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ disappearance signatures in the atmospheric neutrino flux, as illustrated in Fig. 1.
@@
The previous search by IceCube for matter-enhanced signatures used seven years of high-energy atmospheric neutrino data [19, 20], with over 300,000 up-going [^1] We refer to events as up-going when they are reconstructed with $\cos\theta_{z}^{\nu}<0$, indicating that they originated below the horizon. muon tracks in the energy range 0.5 to 10 TeV. The best-fit point was consistent with the standard three-neutrino hypothesis at a p-value of 8%. In addition, no evidence of sterile-driven $\nu_{\mu}$ disappearance was observed in a low-energy dataset using the DeepCore subarray [34], predominantly composed of sub-100 GeV neutrinos.
@@
This Letter provides an updated and improved analysis searching for sterile neutrinos using 11 years of high-energy neutrino data recorded by IceCube. We implement a new event selection that yields a data sample that is larger and higher purity than what was used in previous analyses. After implementing this new selection, the dataset consists of 368,071 up-going $\nu_{\mu}$ events with reconstructed energies ranging from 0.5 to 100 TeV. We separate events with vertices inside or outside the instrumented volume, referred to as “starting” and “through-going events”. This differentiation, coupled with a new energy estimator, enables a more precise characterization of the detector response to $\nu_{\mu}$ interactions, which are substantively different for starting and through-going events. The extended energy range of this analysis requires a significant revision of the systematic uncertainty treatment concerning atmospheric and astrophysical neutrino fluxes. The differentiation between starting and through-going events also introduces a new approach to evaluating systematic uncertainties associated with the glacial ice.
@@
Event selection and reconstruction.—At TeV energies, a muon produced in a $\nu_{\mu}$ [^2] Here, $\nu_{x}$ refers to either neutrinos or antineutrinos of $x$ flavor, with a similar language for their charged-lepton counterparts charged-current (CC) interaction can travel several kilometers in ice [36, 37], emitting Cherenkov radiation and producing an elongated light pattern, referred to as track-like signature, as it traverses the instrumented volume. While a similar pattern is generated by muons originating from cosmic rays in the atmosphere, the effect of Earth’s shielding precludes up-going tracks from this component. Consequently, we can effectively filter out the majority of atmospheric muons by exclusively considering up-going tracks. The track direction is determined with a maximum likelihood method which uses the arrival time distribution of Cherenkov photons registered by the DOMs [38, 39].
@@
The reconstructed $L/E$ distributions for the selected events are shown in Fig. 2, showing a good agreement between data and the prediction at the best-fit point of the analysis in both samples.
@@
Analysis.—Starting and through-going events are binned separately based on their reconstructed energy and the cosine of the zenith angle, $E_{\mathrm{reco}}$ and $\cos\theta_{z}^{\mathrm{reco}}$, respectively. Both datasets use the following binning scheme: 24 bins spanning from 0.5 to 100 TeV in $\log_{10}(E_{\mathrm{reco}})$ and 20 bins spanning from -1 to 0 in $\cos\theta_{z}^{\mathrm{reco}}$. Each bin’s expected and observed event counts are compared using the effective likelihood [45], which is a modified version of the Poisson likelihood to accommodate Monte Carlo statistical uncertainty. Systematic errors are included by weighting the Monte Carlo expectations following the nuisance parameters. A total of 36 parameters are used in this analysis, as detailed below, for which penalty terms are included in the likelihood based on their prior. The final likelihood is given by the product of the effective likelihood and the prior likelihood, $\mathcal{L}=\prod_{i=1}^{N_{bins}}\mathcal{L}_{\mathrm{eff}}(\mu_{i}(\vec{\theta},\vec{\eta});n_{i})\prod_{j=1}^{N_{syst}}G(\eta_{j})$, where $n_{i}$ and $\mu_{i}(\vec{\theta},\vec{\eta})$ are the observed and expected number of events in the i-th bin; and $\vec{\theta}$, $\vec{\eta}$ are the set of physics and nuisance parameters.
@@
The modeling of the conventional flux and its associated uncertainties differ from the previous analysis [^3] The hadronic interactions and cosmic-ray spectrum were modeled with Sibyll2.3c [58] and HillasGaisser H3a [62] in the previous analysis [20]. Systematic uncertainties were accounted for following the Barr prescription [61].. The most notable disparity is evident in the $\nu_{\mu}/\bar{\nu}_{\mu}$ ratio and its associated error, where DAEMONFLUX confines it to $<10\%$ across the entire energy spectrum. For the flux, the error remains below 10% up to 1 TeV, increasing to 30% at higher energies.
@@
Non-conventional neutrino flux: With the maximum reconstructed energy extended from 10 to 100 TeV compared to the previous analysis [20], a more conservative approach has been adopted for the high-energy flux, consisting of astrophysical and prompt components. This analysis represents these fluxes by a broken power law, with uncorrelated Gaussian priors assigned to the normalization and two spectral indices. The prior for the normalization at 100 TeV and both spectral indices are centered at $0.787\times 10^{-18}$ GeV-1sr-1s-1cm-2 and $-2.5$, respectively, and the width of the priors is 46% and 36%. A uniform prior is assumed for the energy pivot point. These priors were chosen to encompass all IceCube astrophysical neutrino flux measurements to date. Mismodeling tests incorporated a galactic component [63], changing the prompt flux and the $\nu/\bar{\nu}$ astrophysical ratio. None of them showed a spurious signal of sterile neutrino in this analysis.
@@
To avoid reliance on Wilk’s theorem, the p-value was alternatively derived by fitting 200 pseudoexperiments, resulting in p$=$3%. Moreover, a Bayesian analysis has been conducted (further details available in [35]), and its results align with those obtained through the frequentist approach.
@@
Figure 3 shows the 90%, 95%, and 99% C.L. contours calculated according to Wilks’ theorem. The sensitivity of this analysis was derived using 500 pseudo-experiments, each of which was generated assuming no sterile neutrino model with the nuisance parameters at their central values. The 90% C.L. preferred region of this analysis is consistent with previous results from IceCube [18, 19, 34].
Figure 3: Frequentist analysis. The 90%, 95%, and 99% C.L. contours, assuming Wilks’ theorem, are shown as dotted, dashed, and solid blue lines, respectively. The red bands show the region where 68% and 95% of the pseudoexperiment 99% C.L. observations lie; the red line corresponds to the median. Previous measurements from IceCube [18, 19, 20, 34] at 90% C.L. are shown in grey.
@@
Conclusions.—Studying the atmospheric $\nu_{\mu}$ spectrum in the TeV regime has evolved into a well-established method for probing sterile neutrinos. This work analyzes 11 years of data collected with the IceCube experiment, incorporating significant enhancements in event selection, reconstruction, and systematic uncertainty treatment compared to earlier studies. The data is consistent with the absence of a sterile neutrino with a 3.1% probability and agrees with previous IceCube analyses. Notably, this analysis introduces for the first time the examination of two distinct event morphologies, namely starting and through-going events, yielding mutually compatible results.

### v2

@@ Abstract
Stringent constraints on sterile-neutrino-induced $\nu_{\mu}$ disappearance arise from IceCube’s study neutrinos produced in the atmosphere with TeV energies, known as high-energy atmospheric neutrinos [18, 19, 20]. IceCube, a gigaton ice-Cherenkov detector near the geographic South Pole [21], detects neutrinos produced in cosmic-ray air showers. The detector consists of an array of photomultiplier tubes in glass pressure housings called “Digital Optical Modules” (DOMs) [22], deployed in 86 strings. In the presence of a sterile neutrino, high-energy atmospheric neutrinos detected by IceCube experience oscillation effects beyond vacuum-like oscillations [23, 24, 25, 26, 27] including, for eV-scale masses, matter-enhanced resonant effects induced as high-energy antineutrinos cross the Earth’s core [28, 29, 30, 31, 32, 33]. Notably, while standard three-neutrino oscillations are negligible at TeV energies due to their mass splitting, the presence of an eV-scale sterile neutrino yields distinct $\nu_{\mu}$ and $\bar{\nu}_{\mu}$ disappearance signatures in the atmospheric neutrino flux, as illustrated in Fig. 1. A detailed explanation of the different oscillation signatures can be found in Ref. [34].
@@
The previous search by IceCube for matter-enhanced signatures used seven years of high-energy atmospheric neutrino data [19, 20], with over 300,000 up-going [^1] We refer to events as up-going when they are reconstructed with $\cos\theta_{z}^{\nu}<0$, indicating that they originated below the horizon. muon tracks in the energy range 0.5 to 10 TeV. The best-fit point was consistent with the standard three-neutrino hypothesis at a p-value of 8%. In addition, no evidence of sterile-driven $\nu_{\mu}$ disappearance was observed in a low-energy dataset using the DeepCore subarray [35, 36], predominantly composed of sub-100 GeV neutrinos.
@@
This Letter provides an updated and improved analysis searching for sterile neutrinos using 10.7 years of high-energy neutrino data recorded by IceCube from May 13th, 2011 to June 7th, 2022. It sets a limit on $\theta_{24}$ compatible with DeepCore’s results and the previous high-energy search. We implement a new event selection that yields a data sample that is larger and higher purity than what was used in previous analyses. After implementing this new selection, the dataset consists of 368,071 up-going $\nu_{\mu}$ events with reconstructed energies ranging from 0.5 to 100 TeV. We separate events with vertices inside or outside the instrumented volume, referred to as “starting” and “through-going events”. This differentiation, coupled with a new energy estimator, enables a more precise characterization of the detector response to $\nu_{\mu}$ interactions, which are substantively different for starting and through-going events. The extended energy range of this analysis requires a significant revision of the systematic uncertainty treatment concerning atmospheric and astrophysical neutrino fluxes. The differentiation between starting and through-going events also introduces a new approach to evaluating systematic uncertainties associated with the glacial ice.
@@
Event selection and reconstruction.—At TeV energies, a muon produced in a $\nu_{\mu}$ [^2] Here, $\nu_{x}$ refers to either neutrinos or antineutrinos of $x$ flavor, with a similar language for their charged-lepton counterparts charged-current (CC) interaction can travel several kilometers in ice [37, 38], emitting Cherenkov radiation and producing an elongated light pattern, referred to as track-like signature, as it traverses the instrumented volume. While a similar pattern is generated by muons originating from cosmic rays in the atmosphere, the earth’s shielding prevents up-going cosmic ray muons from contaminating this sample. Consequently, we can effectively filter out the majority of atmospheric muons by exclusively considering up-going tracks. The track direction is determined with a maximum likelihood method which uses the arrival time distribution of Cherenkov photons registered by the DOMs [39, 40].
@@
The reconstructed $L/E$ distributions for the selected events are shown in Fig. 2, showing a good agreement between data and the prediction at the best-fit point of the analysis in both samples. Different types of events populate different regions of the $L/E$ distribution. Horizontal events predominantly cluster at low $L/E$ values, while vertically up-going events are more common in the high $L/E$ region. The best-fit expectation for $L/E$ shows a dip at 0.3 km/GeV, primarily due to vacuum oscillations. The predicted excess at low $L/E$ arises because the best-fit expectation prefers a higher normalization compared to the null fit, due to fast oscillations happening at $L/E$ below the oscillation maximum. This excess is canceled out by the resonant disappearance effects at $L/E$ 1km/GeV. For a lower mass splitting than the best-fit point, the oscillation maximum shifts to larger $L/E$ values. Conversely, the disappearance effect moves toward lower $L/E$ values if $\Delta m^{2}_{41}$ is large, reaching a point where the sample would only be sensitive to the fast oscillations.
@@
Analysis.—Starting and through-going events are binned separately based on their reconstructed energy and the cosine of the zenith angle, $E_{\mathrm{reco}}$ and $\cos\theta_{z}^{\mathrm{reco}}$, respectively. Both datasets use the following binning scheme: 24 bins spanning from 0.5 to 100 TeV in $\log_{10}(E_{\mathrm{reco}})$ and 20 bins spanning from -1 to 0 in $\cos\theta_{z}^{\mathrm{reco}}$. Each bin’s expected and observed event counts are compared using the effective likelihood [46], which is a modified version of the Poisson likelihood to accommodate Monte Carlo statistical uncertainty. Systematic errors are included by weighting the Monte Carlo expectations following the nuisance parameters. A total of 36 parameters are used in this analysis, as detailed below, for which penalty terms are included in the likelihood based on their prior. The final likelihood is given by the product of the effective likelihood and the systematic prior probability, $\mathcal{L}=\prod_{i=1}^{N_{bins}}\mathcal{L}_{\mathrm{eff}}(\mu_{i}(\vec{\theta},\vec{\eta});n_{i})\prod_{j=1}^{N_{syst}}G(\eta_{j})$, where $n_{i}$ and $\mu_{i}(\vec{\theta},\vec{\eta})$ are the observed and expected number of events in the i-th bin; and $\vec{\theta}$, $\vec{\eta}$ are the set of physics and nuisance parameters.
@@
The modeling of the conventional flux and its associated uncertainties differ from the previous analysis [^3] The hadronic interactions and cosmic-ray spectrum were modeled with Sibyll2.3c [59] and HillasGaisser H3a [63] in the previous analysis [20]. Systematic uncertainties were accounted for following the Barr prescription [62].. The most notable disparity is evident in the $\nu_{\mu}/\bar{\nu}_{\mu}$ ratio and its associated error, where DAEMONFLUX confines it to $<10\%$ across the entire energy spectrum. For the flux, the error remains below 10% up to 1 TeV, increasing to 30% at higher energies. More details about the new parameterization can be found in Ref. [34].
@@
Non-conventional neutrino flux: With the maximum reconstructed energy extended from 10 to 100 TeV compared to the previous analysis [20], a more conservative approach has been adopted for the high-energy flux, consisting of astrophysical and prompt components. This analysis represents these fluxes by a broken power law, with uncorrelated Gaussian priors assigned to the normalization and two spectral indices. The prior for the normalization at 100 TeV and both spectral indices are centered at $0.787\times 10^{-18}$ GeV-1sr-1s-1cm-2 and $-2.5$, respectively, and the width of the priors is 46% and 36%. A uniform prior is assumed for the energy pivot point. These priors were chosen to encompass all IceCube astrophysical neutrino flux measurements to date.
[…]
@@
To avoid reliance on Wilk’s theorem, the p-value was alternatively derived using the Feldman-Cousins procedure [78]. We fitted 200 pseudoexperiments generated at the null point, resulting in p$=$3%. Moreover, a Bayesian analysis has been conducted (further details available in [34]), and its results align with those obtained through the frequentist approach.
@@
Figure 3 shows the 90%, 95%, and 99% C.L. contours calculated according to Wilks’ theorem. The sensitivity of this analysis was derived using 500 pseudo-experiments, each of which was generated assuming no sterile neutrino model with the nuisance parameters at their central values. The 90% C.L. preferred region of this analysis is consistent with previous results from IceCube [18, 19, 35]. This result is compared with previous measurements of $\nu_{\mu}$ disappearance from other experiments in Fig. 4. The 90% C.L. allowed region from this result indicates an increased tension with the constraints from long-baseline experiments.
Figure 3: Frequentist analysis. The 90%, 95%, and 99% C.L. contours, assuming Wilks’ theorem, are shown as dotted, dashed, and solid blue lines, respectively. The red bands show the region where 68% and 95% of the pseudoexperiment 99% C.L. observations lie; the red line corresponds to the median. Previous measurements from IceCube [18, 19, 20, 35, 36] at 90% C.L. are shown in grey.
Figure 4: Comparison to other experiments. The 90% and 99% C.L. contours (blue lines) of this analysis compared to 90% C.L. contours from MiniBooNE-SciBooNE [79, 80], MINOS [81], CDHS, CCFR [7], and SuperK [82].
@@
Conclusions.—Studying the atmospheric $\nu_{\mu}$ spectrum in the TeV regime has evolved into a well-established method for probing sterile neutrinos. This work analyzes 10.7 years of data collected with the IceCube experiment, incorporating significant enhancements in event selection, reconstruction, and systematic uncertainty treatment compared to earlier studies. The data is consistent with the absence of a sterile neutrino with a 3.1% probability and agrees with previous IceCube analyses. Notably, this analysis introduces for the first time the examination of two distinct event morphologies, namely starting and through-going events, yielding mutually compatible results.